One bar or candy A and two bars of candy B have 775 calories. Two bars of candy A and one bar of candy B contain 779 calories. Find the caloric content of each candy bar.
Answer by alicealc(293)

Let:
A = caloric content on candy A
B = caloric content on candy B

(1).. 1A + 2B = 775
(2).. 2A + 1B = 779
multiply (1) by 2:
2A + 4B = 1550
2A + 1B = 779 (-) (subtract these two equations)
---------------
3B = 771
B = 771/3 = 257
substitute B = 257 into (1)
1A + 2(257) = 775
A + 514 = 775
A = 775 - 514 = 261

So, the caloric content in 1 bar of candy A = 261 calories
and the caloric content in 1 bar of candy B = 257 calories